REFL55 has been installed on the AP table. REFL11 has been moved to make space for a 50% beam splitter. The reflected beam from this splitter is about 30% of the transmitted beam power. The reflected beam goes to REFL11 in the current configuration. The DC levels are 1.2V on REFL 11 and 3.5V on the REFL55.
I redid some of the cabling on the table because the we need to choose the heliax cables such that they end up close to the demod board location. As per the 1Y2 (LSC) rack layout given here, some of the PD signals have to arrive at the top and others at the bottom of the LSC rack.
Currently the PDs are connected as follows:
REFL11 PD --> Heliax (ASDD133) (arriving at the top of LSC rack) --> REFL11 Demod Board
REFL55 PD --> Heliax (REFL166) (arriving at the top of LSC rack) --> AS55 Demod Board
AS55 PD --> Heliax (AS166) (arriving at the top of the LSC rack) --> not connected.
We are waiting for the Minicircuits parts to modify the rest of the demod boards.
The heliax cables arriving at the LSC rack are not yet fixed properly. I hope to get this done with Steve's help today.
The first time I noticed that the MC was not locking was after I had finished switching the RF source installation. Before this change the RF modulation frequency (for MC) was 29.485 MHz as read from the Marconi RF Source. We replaced this with a Wenzel crystal source at 29.491 MHz. This may have changed the loop gain.
Today, I changed the MC alignment to optimise the MC lock. Valera pointed out that this is not a desirable solution since it would shift the beam pointing for all components downstream. However, since we are not sure what was the last stable configuration, we decided to stay with the current settings for now and see the trends of several parameters which would tell us if something is drifting and causing the autolocker to fail.
The MC Auto locker is now working okay. However to obtain lock initially we reduced the loop gain by decreasing the VCO gain. We then increased the gain after the autolocker had locked the MC.
I found that the MC autolocker was OFF. Kiwamu says he turned it off because its slow. Suresh says that he has some feelings that maybe something is wrong. I'll let them describe what they know about the MC in an elog.
I checked the trend of the MC and PMC transmissions for the past 30 days:
Looks like the alignment has been drifitng. PMC was corrected recently by Koji, but the alignment of the input beam to the MC or the MC itself has to be fixed. Has someone been twiddling the MC SUS alignment biases??
Since we've got the PRMI locked we now should be able to do more qualitative measurements.
Here is a task list that we will measure/develop in the PRMI condition.
- Optical gain measurements
- Characterization of control loops
- MICH and PRC calibrations
- Noise budget
- Development of automatic noise budget scripts
- Arm loss measurement
- Shnupp asymmetry measurement
My observation wasn't accurate enough.
The looseness came from the fact that the N-SMA bulk heads were slipping on the black plate.
This is actually what Suresh pointed out (see here). So the thickness of the black plate doesn't matter in this case.
Somehow I was able to tighten the bulk heads using two wrenches and I think they are now tight enough so that the heliax's heads don't move any more.
I found that all the Heliax cables landing on the bottom of 1Y2 were too loose.
The looseness basically comes from the fact the black plate is too thick for the Heliax cable to go all the way. It permits the Heliax's heads to rotate freely.
Here are some details about the PRMI locking done last night.
REFL11 has been installed on the AP table. The RF signal from the RFPD is sent by a heliax cable which has been called ASDD133.
Before the beam goes into the RFPD a HWP and PBS are installed such that we can adjust the amount of light entering to the photo diode.
One thing I didn't like was that I had to introduce a big amount of the light into the PD to get a reasonably big RF signal.
I was trying to look for an RF signal by looking at a spectrum analyzer, then I realized that the RF signal at 11 MHz was quite tiny when the DC_MON was less than 1.5 V.
After I increased the amount of the light up to 1.9 V in DC_MON, which sounds already too much, I then got able to see the 11 MHz signal on the analyzer.
Note that I decreased the amount of the light down to 0.5 V after I finished locking the PRMI.
We should make sure what is going on with the 11 MHz modulation.
First I started locking the MIchelson with AS55. The demodulation phase was already somewhat optimized to the I-signal port.
So I decided not to touch the demodulation phase matrix because it may take some times.
After I eliminated electrical offsets in the digital side, I was easily able to lock the Michelson. The control sign was plus.
Then I started playing with the PRC control too. The demodulation phase in REFL11 looked nearly 45 deg although I didn't carefully measure it.
I made a 45 deg rotational matrix to maximize the I-port signal and tried to lock the PRC. Then immediately I was able to lock PRC as well as MICH.
GAIN_MICH = 100
GAIN_PRC = 100
Also GAIN_PRC = -100 gave a carrier resonant lock.
The control filters are the same in MICH and PRC. I used my favorite filters as usual.
FM1 = 1000 : 10
FM6 = 0.1 : 1
FM7 = 1 : 50
Somehow I frequently failed to engage the boost filters (i.e. FM6 and FM7) it looks offsets in the control path kicks either BS or PRM.
The PRMI has been successfully locked
Details will be posted in the morning.
The returning spot diameter on the qpd ~10 mm. In order to reduce the spot size I moved the f 1145 mm lens toward the PRM ~ 25 cm. The spot size was reduced to ~8 mm, 3200 counts.
I'll try to find an other lens tomorrow.
Atm 1, PRM oplev inward path with 2 lens solution: 14 cm gap between F 1145 and F 1545 mm lenses.
Atm 2, The PRM beam size 3 mm and the beam quality is still bad. The BS path only needed alignment.
The little red all terrain cargo wagon 40" x 18" has just arrived on pneumatic wheels.
Model #29, 200 lbs max load at 26 PSI, minimum age requirement 1.5 years
This is a continuation of this
The low pass filter is finally acceptable, and its Bode graph is below (on a ~3Hz frequency span that shows the cutoff frequency is at 0.1Hz)
Jenne went through all the suspension racks and pushed all the connectors.
After pushing them, we had a quick look at those spectra and found no funny noise spectrum except for C1:PRM-SENSOR_UL.
We then checked connection around the SCSI cables and eventually found the connection between ADC_card_0 and a SCSI was loose.
We put short standoffs on the ADC card so that the screws from the SCSI can nicely reach to the ADC card. Now everything looks fine.
SUS diagnostic is quite useful !
Notice that the C1:SUS-ITMX_SENSOR_UL and C1:SUS-MC3_SENSOR_UL spectra fall as 1/f. Jenne suggested that this might indicate that there is a loose electrical connection.
I was charge with making a Non-Inverting Op Amp Low Pass Feedback circuit for Jenne, which may somehow be integrated into a seismometer project she's working on.
Circuit diagram is attached. Calculations show that R1, R2 and C have the following relationship: if R1=10^n, R2=10^(n+1), C=10^(-n-4). For the particular circuit being modeled by the transfer function, R1=100 Ohm, R2=1k Ohm, and C=1uF.
Attached also is the circuit's Bode plot, showing frequency versus gain and phase, respectively. The frequency versus gain graph is true to what the circuit was calculated to generate: a gain of +20 and a cutoff frequency at 200Hz. Not sure what's going on with the frequency verus phase plot.
Here are the free-swinging spectra for the BS, ETMX, ETMY, ITMX, ITMY, MC1, MC2, MC3, and PRM chambers. Kiwamu left the suspensions free for 5 hours this weekend, starting at Sat Apr 30 00:15:26 2011.
This is GPS time 988 182 941. Quick tip: you can do local to GPS time conversions using lalapps_tconvert, which is a lot like tconvert but with special powers. It is installed on pianosa.
$ lalapps_tconvert Sat Apr 30 00:15:26 2011
I generated these figures with the attached Python script, measure.py.
Also, notice that C1:SUS-ETMY_SENSOR_LR, C1:SUS-ITMY_SENSOR_LL, and C1:SUS-PRM_SENSOR_SIDE are a lot noisier above 10 Hz.
Done. C1:PSL-PMC_PMCTRANSPD was improved from ~0.75 to 0.87.
- PMC alignment (Jenne/Koji)
I temporarily turned off the power to the 1Y2 rack this morning while wiring in the binary output adapter board power (+/- 15V) into the cross connects.
The board is now powered and we can proceed to testing if can actually control the LSC whitening filters.
1) Filled in the C1SUS_BS_OLMATRIX properly so as to make the BS oplev work for Steve.
2) Turned on the ITMX damping. Apparently it had tripped this morning, possibly due to work in the lab area.
3) The ETMX FE controller (c1scx) had ADC timed out and died sometime around 8:30 am. The c1x01 (the IOP on the ETMX computer) was also indicating a FB status error (mismatch in DAQ channels).
The reported error in dmesg on c1iscex was:
[1628690.250002] c1spx: ADC TIMEOUT 0 3541 21 3605
[1628690.250002] c1scx: ADC TIMEOUT 0 3541 21 3605
Just to be safe, I rebuilt the c1x01 and c1scx models, ran ./activateDAQ.py, and used the scripts killc1spx, killc1scx, and killc1x01.
I finally restarted the process with startc1x01, startc1scx, and startc1spx. Everything is currently alive and indicating all green.
I think the installation of the PD DC signals are quite important. What to do
1) Connect the DC signals to the right top whitening board (be aware that there may be the modification of the whitening circuit).
2) Reconfigure the LSC model such that the DC signal is passed to the right channels (modify the left top part of the model)
Daytime tasks :
- PRM & BS oplev (Steve)
- LSC binary outputs (Joe/Jamie)
- installation of the REFL55 RFPD (Suresh/Jamie)
- Adjustment of demodulation phases (Kiwamu)
- Bounce-Roll filters on BS and PRM (Suresh/Joe)
- Suspension diagnostic using the free-swinging spectra (Leo)
- PMC alignment (Jenne/Koji)
REFL55 was modified. The noise level confirmed. The PD is now ready to be installed.
Kevin's measurement report told us that something was wrong with REFL55 PD. The transimpedance looked OK, but the noise level was terrible (equivalent to the shotnoise of 14mA DC current).
Rana and I looked at the circuit, and cleaned up the circuit, by removing unnecessary 11MHz notch, 1k shunt resister, and so on.
I made a quick characterization of the PD.
The transimpedance ws measured as a function of the frequency. The resonance was tuned at 55MHz. The notch was tuned at 110MHz in order to reject the second harmonics. The transimpedance was ~540V/A at 55MHz. (For the calibration, I believed the DC transimpedance of 50V/A and 10000V/A for the DC paths of this PD and #1611, respectively, as well as the RF impedance (700V/A0 of #1611.
Output noise levels were measured with various amount of photocurrent using white light from a light bulb. The measurement was perforemed well above the noise level of the measurement instruments.
The measured output noise levels were converted into the equivalent current noise on the PD. The dark noise level agrees with the shot noise level of 1.5mA (i.e. 22pA/rtHz). In deed, the noise level went up x~1.5 when the photocurrent is ~1.4mA.
Also changing the sign of the PRC control gave me the lock of the carrier resonant condition.
The screenshot above is the time series of the error signals when I was locking the PRMI in the sideband resonant condition (i.e. carrier is non-resonant).
Note that I used REFL11 for the PRC control and AS55 for the MICH control as planed.
Details will be posted in the morning.
Today we will try to lock the PRMI.
The Martian wireless bridge has the ethernet cable inserted in the wrong connector.
It should be inserted to one of the four port. Not in the "INTERNET" connector.
Once the connector has been changed, the martian net as well as the internet became accessible from the laptops.
The countersink gives rise to another problem when we mount the N-type-to-SMA bulkhead adaptor. As we are making a circular hole in the plastic strip (instead of a hole with two flat sections) the adaptor is free to turn when we tighten it with a wrench. We currently hold the smooth circular part on the other side with a gripping pliers and while tightening. If that part disappears into the countersink (as seen in the pics) we will not be able to tighten the adaptor sufficiently and consequently we will also not be able to get the heliax connector to be tight.
A better solution would be to use the 1/4-inch plastic L-angle beam which Steve has used on the AS table. In addition to solving this loose connector problem, the beam is also more rigid than the plastic strip.
Due to this loose connection the RF power at 55 MHz varies from -34 dBm to 3 dBm, depending on the angle of the Heliax's head.
What we should do is to make countersinks on the black plate like this:
- AS55 RFPD
With a help from Jamie the AS55 RFPD was installed.
- 55 MHz demodulation board
The AS55 demod board was installed on 1Y2.
- 3-way combiner
ZFSC-3-13 has been installed. All the RF cables from the source side were connected to the combiner.
- installation of the REFL11 RFPD
- DAQ check for AS55 and REFL11
(to be done in the daytime)
- setup AS55 RFPD
- install a demod board for 55 MHz
- install a 3-way RF combiner on EOM.
- prepare 55 MHz RF source (Marconi or RF source box ?)
I swapped the name of two demodulation boards (AS55 and REFL55).
Now the REFL11 and AS55 demodulation boards are ready to go for the PRMI locking.
The physical labels, which are on the front surface of the boards, are also corrected to avoid a confusion.
Here is the latest RF status.
The files are on https://nodus.ligo.caltech.edu:30889/svn/trunk/suresh/40m_RF_upgrade/.
.....Happy.... Birthday.... to.... Joseph... and... Jamie...Happy....Birthday..... to.... You............sing with us........Happy Birthday.....to you
I had a quick look at PRM optical lever.
The He-Ne beam is still successfully coming out from the chamber and I could guide it to the QPD by using steering mirrors.
But the beam size looks too big for the QPD. We should slide the lens which is standing before the injection to get a moderately smaller beam size at the QPD.
- activation of PRM oplev
Today we will try to lock the PRMI. Here is a plan for it.
- setup REFL11 RFPD
- adjustment of each demodulation phase
- AS55_Q ==> BS
- REFL11_I ==> PRM
There is a useful script for this particular job : shutting down all the suspensions and bringing it back to operation after 5 hrs.
It is called opticshudown, which resides in /cvs/cds/rtcds/caltech/c1/scripts/SUS/.
Also I added this script on the list in the wiki where all the scripts will be listed.
If you find any other useful scripts, please add them on the wiki.
I leave all the suspensions free from the watchdogs for 5 hours from now.
- AC coupling for the comparator circuit of the green locking
In order to relieve the power consumption of the RF buffer, ac coupling circuits have been added.
The ac coupling before the buffer amp helps to relieve the power consumption in the chip.
But because of the distortion of the signal (and the limitation of the bandwidth), the output still has some DC (~0.6V).
Therefore, the output is also AC coupled.
Note that the BW pin of BUF634P should be directly connected to -15V in order to keep the bandwidth of the buffer.
The drawings are also uploaded on the green electronics wiki
As a part of the DRMI preparation,
Please DO NOT touch them.
I will check the spectra and the mechanical resonant frequencies on Monday.
Also I will renew all the input matrices of the local dampings based on these free swinging spectra.
[Joe, Jamie, Suresh]
We have installed the IDE to SCSI adaptor module into the 1X2 rack and have connected the AA filter outputs to it.
We have removed the following cables running between the 1X2 and 1X3 racks.
The long twisted pair ribbon cable which previously carried the ADC signals.
1X2-ASC 6, 1X2-ASC 47, 1X2-ASC 9, 1X2-ASC 8, 1X2-ASC 10, 1X2-ASC 7,
CAB-1X2-LSC 42, CAB 1X2-LSC 56, CAB 1X2-LSC 41, CAB 1X2-LSC 43
1X3-2 ASC 47
We have also removed the following by mistake. We will put them back them on Monday
1X2-LSC 21, 1X2-LSC-20.
We have also removed the ASC QPD cables and moved the QPD cards which were present in the middle Eurocate (#2) to the unused Eurocrate at the bottom position (#3).
The binary input cables at the back of the cards require to be supported so that their weight does not pull them out of the sockets at the back of the crates.
Some of the slots where we plan to plug in Demod boards (the 165 MHz boards) are not currently connected to any binary output on the C1:LSC computer. We need these binary controls for the fitlter modules on the cards.
When we eventually begin to use the 15PDs as planned, then we will occupy 30 ADC channels (I & Q outputs). Currently we have just one ADC card installed on the C1:LSC providing 32 ADC channels. Joe found another 16bit 32 channel ADC card in his stash but we need to get a timing+adaptor board for it. In general we are going to need the third Eurocrate.
A platform for the power supply of the RF Distribution box needs to be built and the power supply needs to be moved into the 1X2 rack rather than sit on top of 1X2 rack.
Way back, Jay had D-sub to SCSI adapters made to adapt our existing Sander box AA filters to the new SCSI based IO chassis. However, these did not fit inside the box.
At the time, we simply left the cards outside hanging, which was a hack and needed to be replaced.
Steve modified a black AA filter box so that it could fit the D-sub to SCSI adapter board on it, plus strain relief the SCSI cable, rather than let it hang. The back of the box was cut, and an extending piece of metal attached to the bottom of the box. The adapter board was screwed into the box, the SCSI plugged in, then the SCSI cable is clamped to the extending metal as well.
This modification will be propagated to the 3 remaining AA filter boards using the D-sub to SCSI adapter.
The same modification was carried out at 1X5 for PRM & SRM.
Note: D68L8EX-850Hz are removed and bypassed in 7 channels.
* Farfalla, a lab laptop, seems out of network.
If you look at the real host table instead of the misleading host table in the wiki, you will see that someone has deleted Farfalla from there. She needs to be re-added.
As far as I know, this button works only once after the launch of MEDM...
* reload button on sitemap.adl doesn't work.
First of all, the conclusions / results from the exercise of the Y arm locking yesterday are:
The position of the beam spots on both ETMY and ITMY are now not so bad ( ~ 5 mm off from the center).
The input PZTs are coarsely aligned to the Y arm.
Nevertheless IP_ANG is still too high to come out from the view port at the Y end station.
After the alignments of PRM, SRM and Michelson, POP is still largely clipped.
(what I did)
- Alignments of the Input PZTs
- Adjustment of the demodulation phase for the Y arm PDH.
- Activation of oplev on ITMY
- PDH locking
(Broken or likely broken stuff)
* IP_ANG doesn't give a signal to the digital side.
* ETMY coils look weak and 2 - 3 times weaker than the other test masses. (or OSEM readout gain maybe lower)
* reload button on sitemap.adl doesn't work.
The Y arm has been locked with the IR beam. The purpose is to use the arm as an alignment reference for the input PZTs.
Detail will be posted later. Here is a picture of ITMY suspension. You can see there is a beam spot in the middle of the test mass.
I tried the idea that the PRC can resonate f1 and f2 at the same time if the arm gives the reflection phase to f1 and f2 with the ratio of 1 vs 5.
The details are described on wiki. The point is this removes all of the PRC/SRC/asymmetry mumbo jumbo.
The calculated cavity lengths for f_mod of 11.065399MHz are:
Arm Length: 37.7974 [m]
PRC Length: 6.7538 [m]
SRC Length: 5.39915 [m]
Asymmetry (lx-ly): 0.0342 [m]
Here is the actual values derived from the photos.
Arm Length: 37.54 [m] (0.26m too short)
PRC Length: 6.760 [m] (6mm too long)
SRC Length: 5.415 [m] (16mm too long)
Asymmetry (lx-ly): 0.0266 [m] (8mm too long)
With some assistance from Kiwamu and Koji, I've drawn up the electronics design for the Beat Box for the vertex green locking. The Omingraffle schematic is posted on the Green Locking Wiki page. It's also attached below. Some final touches are necessary before we can Altium this up.
Attachment 1: Schematic of beatbox
Attachment 2: Front and back panel designs.
We are now using the LIGO CDS SVN for storing our control models.
The SVN is at:
The models are under cds_user_apps, then trunk, then approriate subsystem (ISC for c1lsc for example), c1 (for caltech 40m), then models.
We have checked out /cds_user_apps to /opt/rtcds/.
So to find the c1lsc.mdl model, you would go to /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl
This SVN is shared by many people LIGO, so please follow good SVN practice. Remember to update models ("svn update") before doing commits. Also, after making changes please do an update to the SVN so we have a record of the changes.
We are creating soft links in the /opt/rtcds/caltech/c1/core/advLigoRTS/src/epics/simLink/ to the models that you need to build. So if you want to add a new model, please add it to the cds_users_apps SVN in the correct place and create a soft link to the simLink directory.
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1sus.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sus.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1sup.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1sup.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1spy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spy.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1spx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1spx.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1scy.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scy.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1scx.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1scx.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1mcs.mdl -> /opt/rtcds/cds_user_apps/trunk/SUS/c1/models/c1mcs.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1x05.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x05.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1x04.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x04.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1x03.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x03.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1x02.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x02.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1x01.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1x01.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1rfm.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1rfm.mdl
lrwxrwxrwx 1 controls controls 55 Apr 28 14:41 c1dafi.mdl -> /opt/rtcds/cds_user_apps/trunk/CDS/c1/models/c1dafi.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1pem.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1pem.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1mcp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1mcp.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1lsp.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsp.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1lsc.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1lsc.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1ioo.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ioo.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1gpv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gpv.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1gfd.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gfd.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1gcv.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1gcv.mdl
lrwxrwxrwx 1 controls controls 54 Apr 28 14:41 c1ass.mdl -> /opt/rtcds/cds_user_apps/trunk/ISC/c1/models/c1ass.mdl
Since Suresh has installed the RF source box and changed the cable configuration somewhat,
the demodulation phase for the MC locking became off by about 10 degree.
I changed the length of some cables and obtained a good demodulation phase by the same technique as Suresh and Koji did before (see here for detail).
I maximized the Q signal. The lock of the MC looks healthy.
RF Source box has been mounted in the 1X2 rack.
First, I disabled front end starts on boot up, and brought c1sus up. I rebuilt the models for the c1sus computer so they had a new specific_cpu numbers, making the assumption that 0-1 were one real core, 2-3 were another, etc.
Then I ran the startc1SYS scripts one by one to bring up the models. Upon just loading the c1x02 on "core 2" (the IOP), I saw it fluctuate from about 5 to 12. After bringing up c1sus on "core 3", I saw the IOP settle down to about 7 consistently. Prior to hyper-threading it was generally 5.
Unfortunately, the c1sus model was between 60 and 70 microseconds, and was producing error messages a few times a second
[ 1052.876368] c1sus: cycle 14432 time 65; adcWait 0; write1 0; write2 0; longest write2 0
[ 1052.936698] c1sus: cycle 15421 time 74; adcWait 0; write1 0; write2 0; longest write2 0
Bringing up the rest of the models (c1mcs on 4, c1rfm on 5, and c1pem on 6), saw c1mcs occasionally jumping above the 60 microsecond line, perhaps once a minute. It was generally hovering around 45 microseconds. Prior to hyper-threading it was around 25-28 microseconds.
c1rfm was rock solid at 38, which it was prior to hyper-threading. This is most likely due to the fact it has almost no calculation and only RFM reads slowing it down.
c1pem continued to use negligible time, 3 microseconds out of its 480.
I tried moving c1sus to core 8 from core 3, which seemed to bring it to the 58 to 65 microsecond range, with long cycles every few seconds.
I built 5 dummy models (dua on 7, dub on 9, duc on 10, dud on 11, due on 1) to ensure that each virtual core had a model on it, to see if it helped with stabilizing things. The models were basically copies of the c1pem model.
Interestingly, c1mcs seemed to get somewhat better and only taking to 30-32 microseconds, although still not as good as its pre-hyper-threading 25-28. Over the course of several minutes it was no longer having a long cycle.
c1sus got worse again, and was running long cycles 4-5 times a second.
At this point, without surgery on which models are controlling which optics (i.e. splitting the c1sus model up) I am not able to have hyper-threading on and have things working. I am proceeding to revert the control models and c1sus computer to the hyper-threading state.
1) Turn off MC1, MC2, MC3, BS, ITMX, ITMY, PRM, SRM watchdogs.
2) Turn c1sus computer off (sudo shutdown now)
3) Go connect monitor and keyboard to c1sus. Turn c1sus on.
4) Hit "del" key at the right time to go to setup (BIOS).
5) Go to BIOS advanced tab, CPU options, enable Multi-threading.
6) Hit F10 to save and let the computer continue booting.
Once c1sus was up, I noticed several red lights and dead keep alives for the c1sus models.
Typing dmesg on c1sus revealed many messages like:
[ 107.583420] c1x02: cycle 33737 time 20; adcWait 10; write1 0; write2 0; longest write2 0
[ 107.583771] c1x02: cycle 33760 time 19; adcWait 11; write1 0; write2 0; longest write2 0
This indicates the Input/Output Processor (IOP) is not completing its duties within the 15 microseconds (1/64 kHz) that it has. These lines indicate its take 20 or 19 microseconds. (I saw messages ranging from 16 to 22 microseconds).
So this seems to agree with Rolf's observations that hyperthreading can cause a 5-10 microsecond increase in computation time.
So the next thing to do is modify which core the codes are running on, and try to get them paired up on the same physical core.
RF Distribution box has been mounted in the 1Y2 rack and is ready for use.
The box receives 11 and 55 MHz Demod Signals from the RF source located in the 1X2 rack.
Heliax cables have been directly attached to the box and anchored on the side of the 1X2 rack. Here is a list of Helix cables which have been connected so far.
On the back side of 1Y2 rack I found a cable, CAB-1X2-LSC_7, which is supposed to be connected to the whitening filter was disconnected.
I plugged it back and confirmed that the whitening filter is under control of EPICS.
Now all the gain sliders seem to be working because I can change the amplitude of signals with the sliders.
To check if the gain sliders are working or not, I intentionally disconnected all the inputs to the whitening filter.
Then I brought a gain slider of interest to the maximum. Due to the big gain I was easily able to see noise lying above ADC noise.
Also if the gain slider is 0 dB, which is the minimum value, the spectrum becomes just ADC noise.
In this way I checked all the gain sliders from PD1 to PD4. The picture below is just an example screenshot when I was doing this test.
Note that each filer is designed to have two poles at 150 Hz and two zeros at 15 Hz.
While checking whitening filters on the LSC rack, I found some epics controls for the whitening looked not working.
So I powered two crates off : the top one and the bottom one on 1Y3 rack.
These crates contain c1iscaux and c1iscaux2. Then powered them on. But it didn't solve the issue.
I worked on AS_11 today. Its ready for its noise / optical gain calibrations. I have left it on Suresh's desk.
This was one of the 24.5 MHz Black Box (Ben Abbott) style RFPDs rescued from LLO. The tunable inductor that was installed was too small to get the frequency down to 11 MHz and so I swapped in one of the shielded, ferrite core ones from our '7mm' CoilCraft kit. It had a range of 1.2 - 1.8 uH according to the datasheet.
I wasn't able to simulataneously get the peak at 11.06 MHz and the notch at 59.3 MHz and so I took Koji's advice and tuned the peak best. The plot above shows how the notch is slightly off. I think its not a problem; to get it better we would have to change out the inductor for the "2-omega" notch, but I was too lazy. The thinking is that its more important to have the gain be symmetric around the signal readout frequency so as to not imbalance the audio sidebands.
Since this one is going to be AS_11, we think that the 22 MHz signal will be tiny: the transmission of the 11 MHz sidebands to the dark port is small. If we later want to put in a 22 MHz notch anyway, there is space to do this via the 'active notch' pads around the MAX4107.
For the above plot, I used the Jenne laser. The DC output of the PD was ~30 mV (~0.6 mA). The RF drive to the laser was -10 dBm: no saturations. I have calibrated out the cable responses, but not using the 1811 setup, so the absolute calibration has yet to be done.
Also, it needs some new stickers. It would be handy if someone can figure out how to get some sheets of stickers that we can put into the printer. Then we can laser printer all of the data onto the stickers and stick them to the RFPD box.
POX11 (see this entry) is now listed as REFL11 (on the very top row).
We will rename POY11 to POP11 for DRMI locking.
Figure.1 I-Q relative phase measurement as a function of LO power.
Blue curve : relative phase of AS55 that I have modified today (#4572).
Red curve : relative phase of AS11 that I had modified a week ago (#4554). Just for comparison.
The relative phase of AS55 agrees approximately what we expected according to the datasheet of PSCQ-2-51W. We expected 85 degree.
Figure.1 I-Q amplitude imbalance as a function of LO power.
From - 5 dBm to 5 dBm in LO power the imbalance is within 3 %.
But the precision of the measurement is also about 2 % (because I used an oscilloscope). Even so the imbalance is still good.
Some plots will be posted later.